# Looking into the Jet Cone of the Neutrino-Associated Very High Energy Blazar PKS 1424+240

### v1

@@ Abstract
The acceleration process of massive particles as well as the production of very high energy (VHE) photons and neutrinos remains a fundamental challenge in astrophysics. We investigate the parsec-scale properties of the blazar PKS 1424+240, that has been selected on the basis of both gamma-ray and neutrino VHE radiation. We analyze VLBA observations of this BL Lac object, stacking 42 polarization-sensitive images collected in 2009–2025 to enhance the signal and reveal persistent parsec-scale structure. Our observations indicate that this object is viewed inside the jet cone, very close to the axis of its relativistic jet, with a viewing angle of $<0.6^{\circ}$. This effectively maximizes Doppler boosting to values $\sim 30$ and enhances both electromagnetic and neutrino emission in the observer’s direction. Based on polarimetric observations, we unambiguously detect a net toroidal component of the jet’s magnetic field. Blazars with very small jet viewing angles offer a solution to the longstanding mismatch between Doppler factors inferred from VLBI and those derived from VHE observations — the so-called ‘Doppler factor crisis’. We show that relativistic beaming plays the critical role in the gamma-ray and neutrino emission of blazars, with direct implications for models of their multi-messenger emission.
@@ 1 Introduction
Understanding how the universe produces photons and neutrinos at the highest observable energies is of broad interest, as it directly relates to fundamental physics that can only be explored through astronomy. Active galactic nuclei (AGN) are strong emitters of very high energy (VHE) photons (e.g., J. Albert et al., 2007; F. Aharonian et al., 2007) and have been proposed as efficient proton accelerators (starting with V. Berezinsky, 1977), capable of producing neutrinos at TeV–PeV and beyond. Their observed properties are strongly influenced by relativistic beaming effects that also introduce significant observational biases. For example, flux-density-limited samples of radio Very Long Baseline Interferometry (VLBI) selected AGN are dominated by distant beamed blazars with a median redshift of $z\sim 1$ (L. Y. Petrov & Y. Y. Kovalev, 2025). However, the role of beaming at very high energies ($\gtrsim 100$ GeV) is not firmly established. Indeed, while fast variability of VHE gamma-rays implies high values of the Doppler factor $\delta$, the lack of strong apparent superluminal motions in VLBI observed jets suggests the opposite. This is known as the ‘Doppler factor crisis’ (see, e.g., J. Albert et al., 2007; F. Aharonian et al., 2007; B. G. Piner & P. G. Edwards, 2018). Breakthroughs in this area have been rare over the past 35 years, with the Doppler factor crisis posing a significant barrier to progress. The neutrino emission from distant blazars is expected to be beamed, with proton acceleration occurring preferentially along the jet direction (e.g., IceCube Collaboration et al., 2018; A. V. Plavin et al., 2025). In contrast, weaker AGN may emit neutrinos (R. Abbasi et al., 2022) without any preferred direction.
@@
Figure 1: The IceCube skymap cutout surrounding the radio position of the PKS 1424+240 blazar. The color scale represents the local p-value from the 9-year maximum likelihood analysis performed by R. Abbasi et al. (2022).
@@
With a redshift $z=0.605$ (S. Paiano et al., 2017), the BL Lac object PKS 1424+240 (common name OQ 240 (catalog )) is the most distant blazar detected in quiescent VHE (e.g., J. Aleksić et al., 2014; P. Padovani et al., 2022). The all-sky analysis performed by R. Abbasi et al. (2022) has identified it as the possible source of the second-highest neutrino excess in the predefined source list in the Northern sky (see Figure 1). The best-fit neutrino spectrum for PKS 1424+240 is soft with a spectral index of $\gamma=3.5$ (R. Abbasi et al., 2022). The IceCube signal is dominated by neutrinos with relatively low, $\lesssim 10$ TeV, energies. The time-integrated IceCube analysis is predominantly sensitive to persistent emission. Together with off-flare VHE detections in gamma-rays, this motivates the study of time-independent properties of PKS 1424+240. Here, we address the question, which general property of PKS 1424+240 may be responsible for its exceptional VHE photon and neutrino fluxes.
@@
We use radio VLBI, which is the only technique to date that is able to directly probe the sites of neutrino and gamma-ray production, resolving structures within several parsecs from the central engine. For the redshift of PKS 1424+240, the projected angular scale of 1 mas corresponds to 6.70 pc, assuming the E. Komatsu et al. (2009) cosmological parameters. A quick examination of PKS 1424+240 has revealed that its jet has an exceptionally large apparent opening angle $\varphi_{\mathrm{app}}$. A. B. Pushkarev et al. (2017) found $\varphi_{\mathrm{app}}=65\arcdeg$, among the most extreme values observed for blazars. Geometrically, this indicates a small viewing angle $\theta$ to the line of sight and, consequently, strong Doppler boosting. To further investigate the source, we updated its stacked 15 GHz image by the Very Long Baseline Array (VLBA) using all available data and increasing the number of stacked epochs by about ×$\times 4$. The results of our serendipitous findings are presented in this paper.
2 VLBA polarization imaging
2.1 Long-term 15 GHz MOJAVE observations: image stacking
@@
Stokes I and polarization observation of PKS 1424+240 in the MOJAVE program [^1] https://www.cv.nrao.edu/MOJAVE/ (MOJAVE: Monitoring Of Jets in Active galactic nuclei with VLBA Experiments, see M. L. Lister et al., 2018, and references therein) began in May 2009 and we report results through January 2025. To increase the dynamic range of a restored map, a set of single-epoch images is combined and averaged, i.e., stacked, by aligning them using the VLBI core position derived from visibility model fitting (e.g., M. L. Lister et al., 2021). Stacking not only improves the sensitivity and dynamic range but also effectively reconstructs more complete source morphology, particularly in the low-brightness areas. It can also reveal the whole jet channel, not just regions of enhanced radiation at a given epoch. It is beneficial for studies of jets in Stokes I and polarized emission (e.g. A. B. Pushkarev et al., 2017, 2023; Y. Y. Kovalev et al., 2020; D. I. Zobnina et al., 2023).
@@
We made stacked maps of PKS 1424+240 in Stokes I, Q, and U (Figure 2), using 42 available epochs that span approximately 15 years, following a procedure described in A. B. Pushkarev et al. (2023). For deep CLEANing in Stokes I, Q, and U, we applied the algorithm with the residual entropy-based stopping criterion developed by D. C. Homan et al. (2024). The linear polarization image has been corrected for Ricean bias (J. F. C. Wardle & P. P. Kronberg, 1974). We also applied a debiasing procedure to the CLEAN images to effectively suppress the sidelobe artifacts from deconvolution (Appendix B in A. B. Pushkarev et al., 2023).
@@
Figure 2: VLBA stacked image for the blazar PKS 1424+240 at 15 GHz. Stokes I is shown by contours, the first contour corresponds to 3×3$\times$3 × the image rms level. Linear polarization intensity is presented by color, directions of electric vector position angle (EVPA) — by sticks. The observation dates of the 42 images used in the stacked image are indicated in the inset. A circular restoring beam is shown at the full width at half maximum (FWHM, 0.8 mas) level in the bottom left corner. FITS files of the stacked Stokes I, Q, and U images are published as electronic supplementary material. We also supply an animation of Stokes I cumulative stacking.
@@
This debiasing procedure has significantly reduced imaging artifacts, and we argue that the weak structures observed around the core of PKS 1424+240 in the stacked images (Figure 2) represent genuine reconstructed jet emission. The northern declination of PKS 1424+240 ensures robust VLBA $uv$-coverage with no significant gaps. The weak polarized emission stacks coherently over multiple epochs, and is not expected to result from residual polarization leakage effects. The reduced levels of detectable 15 GHz Stokes I emission in the North-West direction next to the bright core component possibly results from residual effects of self-calibration which only included emission to the South-East in the self-calibration model in most individual epochs.
In order to check for consistency of the stacked 15 GHz image, we have processed archival 1.6 GHz VLBA data for PKS 1424+240, experiment BS257A, epoch 9 May 2017. At ten times lower frequency, the VLBA is substantially more sensitive to distant optically thin jet emission and probes angular scales ten times larger than our MOJAVE stacked image. With a total on-source time of about 3 hr, these single epoch low-frequency observations reach a high dynamic range of $\sim 7,000:1$ and confirm the emission around the VLBI core, including the North-West direction (Figure 4). We note that our stacked image at 24 GHz also detects this emission (Figure 5).
2.2 Rotation Measure at 15, 24, and 43 GHz
As RM data is not available for all 42 epochs used in Figure 2, and individual epochs do not have RM data for the full jet region, we cannot apply a precise correction for the Faraday effect on EVPA. The median magnitude of RM for PKS 1424+240 corresponds to a EVPA rotation of $\sim 6^{\circ}$, which is comparable to our EVPA errors from calibration and imaging.
The low magnitude and variability of RM especially close to the core of PKS 1424+240 are indicative of either a weak magnetic field or low thermal electron density along the line-of-sight. We may be seeing the ‘sweeping away’ of the magneto-ionic medium, this would result in a reduction in the magnitude of RM. This effect would be more prominent within the center of the jet, while the edges would be less affected, which is supported by the magnitude and variability of RM being typical of other sources for the jet region of PKS 1424+240.
@@ 3.1 Observational signatures and magnetic field
We have collected striking observational signatures of a jet that is viewed inside its cone. The stacked Stokes I and linear polarization images reveal a structure in the southeast direction, while significant emission is also present in all other directions around the core region in Stokes I and linear polarization (Figure 2). This is extended even to deprojected kiloparsec scales, see Figure 4. Furthermore, the stacked polarization image exhibits a remarkably uniform distribution of EVPAs; it presents a pattern of “diverging rays”, which start from the core. The low rotation measure and its minimal variation suggest that the jet is propagating toward the observer, reducing the density of the interstellar medium along its path. The low jet speeds, observed by VLBI (subsection 3.2), are consistent with a small jet viewing angle, $\theta$.
We reconstruct the magnetic field structure, implementing a $90^{\circ}$ EVPA rotation across the entire map. This assumes that RM as well as relativistic effects do not affect the observed projected magnetic field direction significantly, see subsection 2.2 and M. Lyutikov et al. (2005). The results are shown in Figure 3, which (after J. R. R. Tolkien) we call ‘the Eye of Sauron’ due to its striking nature. Linear Integral Convolution (LIC; B. Cabral & L. C. Leedom, 1993) is applied to the polarized intensity to visualize the morphology of magnetic field lines by locally blurring intensity variations along the magnetic field direction, effectively tracing its structure. The image reveals a significant toroidal magnetic field component of a current-carrying jet, which flows almost directly toward Earth. This is consistent with a widely discussed overall helical magnetic field structure (R. Blandford et al., 2019).
@@ 3.2 Basic jet parameters
M. L. Lister et al. (2021) have measured apparent kinematics of two robust components in the emission of PKS 1424+240. Their apparent speed, measured in units of the speed of light, is $\beta_{\mathrm{app,1}}=2.83\pm 0.89$ and $\beta_{\mathrm{app,2}}=1.91\pm 0.18$. These moderate values are consistent with other measurements for the same source (e.g., P. Padovani et al., 2022; E. Kun & A. Medveczky, 2023). A median intrinsic full opening angle of the jet was estimated by A. B. Pushkarev et al. (2017) to be $\alpha_{\mathrm{int}}=1.2^{\circ}$ for Fermi-detected AGN in MOJAVE sample. We assume this value for PKS 1424+240. Looking into the jet cone condition is defined by the requirement for the jet viewing angle $\theta<\alpha_{\mathrm{int}}/2$. This requires the Doppler factor to reach its maximum possible value of twice the Lorentz factor, $\delta\approx 2\Gamma$.
@@
We apply a simplified assumption that the jet is observed at a half of its half opening angle, $\theta=\alpha_{\mathrm{int}}/4$. This is in agreement with simulations by Pashchenko et al. (in prep.). We also assume a relation between $\alpha_{\mathrm{int}}$ and $\Gamma$, as deduced observationally by A. B. Pushkarev et al. (2017) and predicted by hydrodynamical and magnetic acceleration models of relativistic jets (R. D. Blandford & A. Königl, 1979; S. S. Komissarov et al., 2007; E. Clausen-Brown et al., 2013). As a result, we find that the following set of parameters for a basic relativistic jet model are consistent with the above values and considerations: $\Gamma=16$ (plasma flow speed $\beta=0.998$), $\alpha_{\mathrm{int}}=1.2^{\circ}$, $\theta=0.3^{\circ}$, $\beta_{\mathrm{app}}=2.8$, and Doppler factor $\delta=32$. The Lorentz factor is within its typical values from MOJAVE population modeling (M. L. Lister et al., 2019), as expected from the chosen conservative parameter assumptions. The Doppler factor is three times higher than the median values for the MOJAVE sample (M. L. Lister et al., 2021; D. C. Homan et al., 2021), which agrees with recent finding of A. V. Plavin et al. (2025) for neutrino-selected blazars. Note that $\Gamma$ and $\delta$ estimates are not highly sensitive to the exact value of $\theta$, as long as the line of sight remains within the jet cone. Assuming a jet viewing angle in the range $0.1^{\circ}<\theta<1.0^{\circ}$ and using the measured $\beta_{\mathrm{app}}$ values, the Doppler factors lie within the range $1652>\delta>$ 16. Although the geometric probability of a near-zero viewing angle is low, the fraction of such jets in VLBI flux density-selected samples has been shown to be on the order of a few percent due to Doppler bias (M. L. Lister et al., 2019).
@@
Due to strong projection effects and synchrotron opacity, the jet is expected to become visible many parsecs downstream from its true base. As a result, its observed radio core appears less active and variable than typically expected for highly boosted blazars. Moreover, the core and jet emission are superimposed, resulting in a significant underestimation of the intrinsic core brightness temperature, which in turn leads to an incorrect estimation of jet parameters in D. C. Homan et al. (2021).

### v2

@@ Abstract
Context. The acceleration process of massive particles as well as the production of very high energy (VHE) photons and neutrinos remains a fundamental challenge in astrophysics.
Aims. We investigate the parsec-scale jet structure and magnetic field of the blazar PKS 1424+240, that has been selected on the basis of strong VHE gamma-ray emission and identified with one of the highest peaks in the IceCube 9-year neutrino sky.
Methods. We analyze 15 GHz VLBA observations of this BL Lac object, stacking 42 polarization-sensitive images collected in 2009–2025 to enhance the signal and reveal persistent parsec-scale structure.
Results. Our observations uncover a rare scenario. The object is viewed inside the jet cone, very close to the axis of its relativistic jet, with a viewing angle of $<0.6^{\circ}$. This effectively maximizes Doppler boosting to values $\sim 30$ and enhances both electromagnetic and neutrino emission in the observer’s direction. Based on polarimetric observations, we unambiguously detect a net toroidal component of the jet’s magnetic field, indicating a current carrying jet flowing almost directly towards our line of sight.
Conclusions. Blazars with very small jet viewing angles offer a solution to the longstanding mismatch between Doppler factors inferred from low VLBI apparent jet speed and those derived from VHE observations — the so-called “Doppler factor crisis”. We show that relativistic beaming plays the critical role in the gamma-ray and neutrino emission of blazars, with direct implications for models of their multi-messenger emission.
Key Words.:
Neutrinos – Radio continuum: galaxies – Galaxies: jets – BL Lacertae objects: individual: PKS 1424+240
[…]
@@ 1 Introduction
Understanding how the universe produces photons and neutrinos at the highest observable energies is of broad interest, as it directly relates to fundamental physics that can only be explored through astronomy. Active galactic nuclei (AGN) are strong emitters of very high energy (VHE) photons (e.g., Albert et al., 2007; Aharonian et al., 2007) and have been proposed as efficient proton accelerators (starting with Berezinsky, 1977), capable of producing neutrinos at TeV–PeV and beyond. Their observed properties are strongly influenced by relativistic beaming effects that also introduce significant observational biases. For example, flux-density-limited samples of radio Very Long Baseline Interferometry (VLBI) selected AGN are dominated by distant beamed blazars with a median redshift of $z\sim 1$ (Petrov & Kovalev, 2025). However, the role of beaming at very high energies ($\gtrsim 100$ GeV) is not firmly established. Indeed, while fast variability of VHE gamma-rays implies high values of the Doppler factor $\delta$, the lack of strong apparent superluminal motions in VLBI observed jets suggests the opposite. This is known as the “Doppler factor crisis” (see, e.g., Albert et al., 2007; Aharonian et al., 2007; Piner & Edwards, 2018). Breakthroughs in this area have been rare over the past 35 years, with the Doppler factor crisis posing a significant barrier to progress. The neutrino emission from distant blazars is expected to be beamed, with proton acceleration occurring preferentially along the jet direction (e.g., IceCube Collaboration et al., 2018; Plavin et al., 2025). In contrast, weaker AGN may emit neutrinos (Abbasi et al., 2022) without any preferred direction.
@@
Figure 1: The IceCube skymap cutout surrounding the radio position of the PKS 1424+240 blazar. The color scale represents the local p-value from the 9-year maximum likelihood analysis performed by Abbasi et al. (2022).
@@
With a redshift $z=0.605$ (Paiano et al., 2017), the BL Lac object PKS 1424+240 (common name OQ 240) is the most distant blazar detected in quiescent VHE (e.g., Aleksić et al., 2014; Padovani et al., 2022). The all-sky analysis performed by Abbasi et al. (2022) has identified it as the possible source of the second-highest neutrino excess in the predefined source list in the Northern sky (see Figure 1). The best-fit neutrino spectrum for PKS 1424+240 is soft with a spectral index of $\gamma=3.5$ (Abbasi et al., 2022). The IceCube signal is dominated by neutrinos with relatively low, $\lesssim 10$ TeV, energies. The time-integrated IceCube analysis is predominantly sensitive to persistent emission. Together with off-flare VHE detections in gamma-rays, this motivates the study of time-independent properties of PKS 1424+240. Here, we address the question, which general property of PKS 1424+240 may be responsible for its exceptional VHE photon and neutrino fluxes.
@@
We use radio VLBI, which is the only technique to date that is able to directly probe the sites of neutrino and gamma-ray production, resolving structures within several parsecs from the central engine. For the redshift of PKS 1424+240, the projected angular scale of 1 mas corresponds to 6.70 pc, assuming the Komatsu et al. (2009) cosmological parameters. A quick examination of PKS 1424+240 has revealed that its jet has an exceptionally large apparent opening angle $\varphi_{\mathrm{app}}$. Pushkarev et al. (2017) found $\varphi_{\mathrm{app}}=65\degr$, among the most extreme values observed for blazars. Geometrically, this indicates a small viewing angle $\theta$ to the line of sight and, consequently, strong Doppler boosting. To further investigate the source, we updated its stacked 15 GHz image by the Very Long Baseline Array (VLBA) using all available data and increasing the number of stacked epochs by about ×$\times 4$. The results of our serendipitous findings are presented in this paper.
2 15 GHz VLBA polarization imaging
@@
Stokes I and polarization observation of PKS 1424+240 in the MOJAVE program [^1] https://www.cv.nrao.edu/MOJAVE/ (MOJAVE: Monitoring Of Jets in Active galactic nuclei with VLBA Experiments, see Lister et al., 2018, and references therein) began in May 2009 and we report results through January 2025. To increase the dynamic range of a restored map, a set of single-epoch images was combined and averaged, i.e., stacked, by aligning them using the VLBI core position derived from visibility model fitting (e.g., Lister et al., 2021). Stacking not only improves the sensitivity and dynamic range but also effectively reconstructs more complete source morphology, particularly in the low-brightness areas. It can also reveal the whole jet channel, not just regions of enhanced radiation at a given epoch. Image stacking is beneficial for studies of jets in Stokes I and polarized emission (e.g. Pushkarev et al., 2017, 2023; Kovalev et al., 2020; Zobnina et al., 2023).
@@
We made stacked maps of PKS 1424+240 in Stokes I, Q, and U (Figure 2), using 42 available epochs that span approximately 15 years, following a procedure described in Pushkarev et al. (2023). For deep CLEANing in Stokes I, Q, and U, we applied the algorithm with the residual entropy-based stopping criterion developed by Homan et al. (2024). The linear polarization image has been corrected for Ricean bias (Wardle & Kronberg, 1974). We also applied a debiasing procedure to the CLEAN images to effectively suppress the sidelobe artifacts from deconvolution (Appendix B in Pushkarev et al., 2023).
@@
Figure 2: VLBA stacked image for the blazar PKS 1424+240 at 15 GHz. Stokes I is shown by contours, the first contour corresponds to 3×3$\times$3 × the image rms level. Linear polarization intensity is presented by color, directions of electric vector position angle (EVPA) — by sticks. The observation dates of the 42 images used in the stacked image are indicated in the inset. A circular restoring beam is shown at the full width at half maximum (FWHM, 0.8 mas) level in the bottom left corner. FITS files of the stacked Stokes I, Q, and U images are available at the CDS. We also supply an animation of Stokes I cumulative stacking (available online).
@@
This debiasing procedure has significantly reduced imaging artifacts, and we argue that the weak structures observed around the core of PKS 1424+240 in the stacked images (Figure 2) represent genuine reconstructed jet emission. The northern declination of PKS 1424+240 ensures robust VLBA $uv$-coverage with no significant gaps. The weak polarized emission stacks coherently over multiple epochs, and is not expected to result from residual polarization leakage effects. The reduced levels of detectable 15 GHz Stokes I emission in the North-West direction next to the bright core component possibly results from residual effects of self-calibration which only included emission to the South-East in the self-calibration model in most individual epochs. In order to check for consistency the Figure 2 Stokes I image, we have processed archival 1.6 GHz VLBA data for PKS 1424+240, see subsection A.1.
@@ 3.1 Observational signatures and magnetic field
We have collected striking observational signatures of a jet that is viewed inside its cone. The stacked Stokes I and linear polarization images reveal a structure in the southeast direction, while significant emission is also present in all other directions around the core region in Stokes I and linear polarization (Figure 2). This is extended even to deprojected kiloparsec scales, see Figure 4. Furthermore, the stacked polarization image exhibits a remarkably uniform distribution of EVPAs; it presents a pattern of “diverging rays”, which start from the core. Rotation measure 15/24/43 GHz VLBA results are presented in subsection A.2. The low rotation measure and its minimal variation suggest that the jet is propagating toward the observer, reducing the density of the interstellar medium along its path. The low jet speeds, observed by VLBI (subsection 3.2), are consistent with a small jet viewing angle, $\theta$.
We reconstruct the magnetic field structure, implementing a $90^{\circ}$ EVPA rotation across the entire map. This assumes that RM as well as relativistic effects do not affect the observed projected magnetic field direction significantly, see subsection A.2 and Lyutikov et al. (2005). The results are shown in Figure 3, which (after J. R. R. Tolkien) we call ‘the Eye of Sauron’ due to its striking nature. Linear Integral Convolution (LIC; Cabral & Leedom, 1993) is applied to the polarized intensity to visualize the morphology of magnetic field lines by locally blurring intensity variations along the magnetic field direction, effectively tracing its structure. The image reveals a significant toroidal magnetic field component of a current-carrying jet, which flows almost directly toward Earth. This is consistent with a widely discussed overall helical magnetic field structure (Blandford et al., 2019).
@@ 3.2 Basic jet parameters
Lister et al. (2021) have measured apparent kinematics of two robust components in the emission of PKS 1424+240. Their apparent speed, measured in units of the speed of light, is $\beta_{\mathrm{app,1}}=2.83\pm 0.89$ and $\beta_{\mathrm{app,2}}=1.91\pm 0.18$. These moderate values are consistent with other measurements for the same source (e.g., Padovani et al., 2022; Kun & Medveczky, 2023). A median intrinsic full opening angle of the jet was estimated by Pushkarev et al. (2017) to be $\alpha_{\mathrm{int}}=1.2^{\circ}$ for Fermi-detected AGN in MOJAVE sample. We assume this value for PKS 1424+240. Looking into the jet cone condition is defined by the requirement for the jet viewing angle $\theta<\alpha_{\mathrm{int}}/2$. This requires the Doppler factor to reach its maximum possible value of twice the Lorentz factor, $\delta\approx 2\Gamma$.
@@
We apply a simplified assumption that the jet is observed at a half of its half opening angle, $\theta=\alpha_{\mathrm{int}}/4$. This is in agreement with simulations by Pashchenko et al. (in prep.). We also assume a relation between $\alpha_{\mathrm{int}}$ and $\Gamma$, as deduced observationally by Pushkarev et al. (2017) and predicted by hydrodynamical and magnetic acceleration models of relativistic jets (Blandford & Königl, 1979; Komissarov et al., 2007; Clausen-Brown et al., 2013). As a result, we find that the following set of parameters for a basic relativistic jet model are consistent with the above values and considerations: $\Gamma=16$ (plasma flow speed $\beta=0.998$), $\alpha_{\mathrm{int}}=1.2^{\circ}$, $\theta=0.3^{\circ}$, $\beta_{\mathrm{app}}=2.8$, and Doppler factor $\delta=32$. The Lorentz factor is within its typical values from MOJAVE population modeling (Lister et al., 2019), as expected from the chosen conservative parameter assumptions. The Doppler factor is three times higher than the median values for the MOJAVE sample (Lister et al., 2021; Homan et al., 2021), which agrees with recent finding of Plavin et al. (2025) for neutrino-selected blazars. Note that $\Gamma$ and $\delta$ estimates are not highly sensitive to the exact value of $\theta$, as long as the line of sight remains within the jet cone. Assuming a jet viewing angle in the range $0.1^{\circ}<\theta<1.0^{\circ}$ and using the measured $\beta_{\mathrm{app}}$ values, the Doppler factors lie within the range $1652>\delta>$ 16. Although the geometric probability of a near-zero viewing angle is low, the fraction of such jets in VLBI flux density-selected samples has been shown to be on the order of a few percent due to Doppler bias (Lister et al., 2019).
@@
Due to strong projection effects and synchrotron opacity, the jet is expected to become visible many parsecs downstream from its true base. As a result, its observed radio core appears less active and variable than typically expected for highly boosted blazars. Moreover, the core and jet emission are superimposed, resulting in a significant underestimation of the intrinsic core brightness temperature, which in turn leads to an incorrect estimation of jet parameters in Homan et al. (2021). A moderate flare occurred in the core of PKS 1424+240 during 2019–2020; the core flux density has reached 0.27 Jy, while the median value over all epochs is 0.18 Jy. It did not affect these estimates significantly.